# CountingFruit: Real-Time 3D Fruit Counting with Language-Guided Semantic Gaussian Splatting

### v1

@@ 4.3 Instance Counting Accuracy
Figure 3 provides qualitative visualizations of our pipeline on four representative orchard scenes. In this subsection, we evaluate the effectiveness of our full pipeline in terms of counting recall, using both manually annotated trees (Tree01–03) and the full Fuji-SfM benchmark. Table 3 compares our method with three baselines: NeRF [15]+Ours, 3DGS [16]+Ours, and FruitNeRF [9] with its best-performing U-Net-Big-model (U-Net-B) segmentation backbone.
Across all scenes, FruitLangGS achieves the most consistent and accurate instance counts. On Tree03, our method correctly identifies 287 out of 291 fruits, yielding a recall of 98.6%. This is 31.9% higher than the 66.7% recall achieved by 3DGS+Ours, and 7.9% higher than FruitNeRF. On Tree01 and Tree02, our method performs on par with FruitNeRF, with 96.6% and 97.3% recall respectively. In the Fuji-SfM datasets, FruitLangGS detects 1443 out of 1455 ground-truth instances, achieving a recall of 99.2%. This matches the overall performance of FruitNeRF (1459 detections), which reports a slightly inflated recall of 100.3%, likely due to oversegmentation or false-positive clustering around ambiguous fruit boundaries. In contrast, our method maintains recall close to 100% without over-counting, demonstrating its robustness across species, geometries, and occlusion conditions. Also, while 3DGS+Ours offers moderately improved recall due to sharper geometry, it still fails to match the consistency of our adaptive Gaussian reconstruction combined with semantic filtering and clustering.
@@ 4.4 Semantic Embedding Ablation
Table 4 presents an ablation study designed to evaluate the contribution of the language-conditioned semantic embedding module. The full method, referred to as ”with Language-Conditioned Semantic Embedding,” is described in Sec. 3.2. It augments each 3D Gaussian $\mathcal{G}_{i}$ with a compact language feature vector derived from the CLIP model [18] and compressed through a lightweight autoencoder. The first ablation variant disables this semantic embedding by setting the feature level to 0, thereby removing $\mathbf{f}_{i}$ from training and inference. The second variant replaces our embedding module with the Fruit Field representation introduced in FruitNeRF [9], and adopts its original DBSCAN-based clustering strategy.
Experimental results indicate that the semantic embedding plays a critical role in accurate fruit counting. When the embedding is removed, performance degrades substantially. On Tree03, recall drops from 98.6% (287 / 291) to 70.1% (204 / 291), corresponding to an absolute reduction of 28.5%. Tree01 and Tree02 show similar trends, with recall decreasing by 17.8% and 15.0%, respectively. This confirms that semantic guidance is essential for resolving occluded or tightly clustered instances. Across the full Fuji-SfM dataset, recall falls from 99.2% to 81.2% without semantic embedding. Replacing our language-aligned embedding with the Fruit Field approach recovers part of the gap, but still underperforms in all cases. For example, Tree03 improves to 86.3% (251 / 291), which remains 12.3% lower than our method. Similar deficits of 8.3% and 7.9% persist on Tree01 and Tree02, respectively.
@@ 5 Conclusion and Future Work
We present FruitLangGS, a real-time framework for 3D fruit counting that integrates adaptive Gaussian splatting, language-conditioned semantic embedding, and prompt-driven instance estimation. Each module is designed to address specific challenges in orchard-scale perception: efficient rendering for large scenes, open-vocabulary semantic controllability, and accurate clustering under occlusion. Our reconstruction module achieves sub-second rendering by introducing spatial pruning and tile-level scheduling, enabling over 300 FPS. The semantic embedding module encodes multi-level CLIP-aligned features into each 3D Gaussian, supporting language-based filtering without per-frame supervision. The final counting stage leverages distribution-aware sampling and geometric refinement to yield accurate fruit instance estimates. Compared to the state-of-the-art FruitNeRF, FruitLangGS offers improvements in rendering speed (up to 50× faster) and memory usage, while achieving comparable or better counting accuracy. On the Fuji-SfM benchmark, our method reaches 99.2% recall, closely matching FruitNeRF’s 100.3% but without overcounting. Importantly, FruitLangGS generalizes across species and scenes without requiring model retraining for each fruit or tree. In future work, we plan to adapt this framework to embodied AI settings for agricultural simulation and task planning, and extend the semantic vocabulary to support multi-class separation and recognition of other crops and field objects.

### v2

@@ 4.3 Instance Counting Accuracy
Figure 3 shows qualitative results of our pipeline on nine orchard scenes, including both benchmark datasets and newly introduced sequences. These visualizations highlight the robustness of FruitLangGS under diverse lighting, occlusion, and structural configurations. Fruit counting remains the most intuitive and widely adopted evaluation metric in 3D orchard analysis. We evaluate recall against ground-truth annotations where available. Table 3 compares our method with three baselines: NeRF [16]+Ours, 3DGS [17]+Ours, and FruitNeRF [9], evaluated under its best-performing configuration termed U-Net-B. This setting combines U-Net-generated binary fruit masks with the larger FruitNeRF-B backbone, yielding optimal counting performance among the variants reported in the original study.
On the Tree03, our method identifies 287 out of 291 fruits, achieving 98.6% recall, 31.9% higher than 3DGS+Ours and 7.9% higher than FruitNeRF. On Tree01 and Tree02, our method performs comparably to FruitNeRF, achieving 96.6% and 97.3% recall, respectively. Across the large-scale Fuji-SfM set, FruitLangGS detects 1443 of 1455 fruits (99.2% recall), matching FruitNeRF (1459 detections) but without inflated oversegmentation. Furthermore, on Pfuji-Size_Orch2018 and Orch2020 subsets, our method reaches recall rates of 91.8% and 91.3%, outperforming all baselines. For the Galafab-west, Schnico-Red-east, and Schniga-Schnico-west datasets, where per-tree annotations are unavailable, we report estimated counts from full-scene reconstructions. These results affirm the generalizability of FruitLangGS to unseen domains with complex occlusions and variable capture conditions.
@@ 4.4 Semantic Embedding Ablation
Table 4 presents an ablation study to evaluate the role of the language-conditioned semantic embedding module in fruit counting. The full method, described in Sec. 3.2, augments each 3D Gaussian $\mathcal{G}_{i}$ with a language feature $\mathbf{f}_{i}$ extracted from CLIP and compressed through an autoencoder. We compare it against two simplified variants. The first removes semantic embedding entirely by setting $\mathbf{f}_{i}=0$ during training and inference. The second replaces our module with the Fruit Field representation and its DBSCAN-based clustering, as proposed in FruitNeRF [9].
The results show that semantic embedding is essential for accurate counting. Without embedding, recall drops significantly across all datasets. For example, Tree03 recall decreases from 98.6% to 70.1%, and Fuji-SfM drops from 99.2% to 81.2%. Similar trends are observed on Tree01 and Tree02, with absolute reductions of 17.8% and 15.0%, respectively. Replacing our module with Fruit Field improves performance over the embedding-free variant. However, it still lags behind our method. On Tree03, recall reaches 86.3%, which is 12.3% lower than our best result. On Tree01 and Tree02, the gaps remain at 8.3% and 7.9%. These findings confirm that language-guided semantic features are critical for resolving occlusion and separating tightly clustered fruits.
@@ 5 Conclusion and Future Work
FruitLangGS introduces a real-time framework for open-vocabulary 3D fruit counting by coupling adaptive Gaussian rendering with language-guided semantic filtering. This design eliminates the need for category-specific training or per-frame annotations, enabling prompt-driven counting directly on reconstructed scenes. Experiments across nine real-world orchard datasets demonstrate that our method achieves consistent high recall, robust instance separation under occlusion, and rendering speeds, outperforming prior baselines in both accuracy and efficiency. These results validate the core that integrating geometric structure with compact language-aligned embeddings enables scalable, generalizable perception in complex agricultural environments. Future work will explore deploying FruitLangGS in embodied settings, such as robotic arms or mobile platforms, for vision-language-guided manipulation and task execution. We also plan to accelerate training via optimized splatting strategies, further enhancing the framework’s readiness for real-time, on-device agricultural applications.